OpenStudy (anonymous):

The life expectancy of a typical lightbulb is normally distributed with a mean of 2,000 hours and a standard deviation of 27 hours. What is the probability that a lightbulb will last between 1,975 and 2,050 hours?

OpenStudy (anonymous):

is D correct

OpenStudy (cj49):

P(1,975≤X≤2,050)=Φ(2,050−2,00027)−Φ(1,975−2,00027) Where Φ(x) is the CDF for the Normal Distribution.

OpenStudy (cj49):

P(1,975≤X≤2,050)=Φ(2,050−2,000/27)−Φ(1,975−2,000/27)=0.968−0.1772

OpenStudy (cj49):

yes D is d ans